### v1

@@ 3.3 Dynamic Mixture-of-Experts
$\mathbf{G}=\operatorname{Softmax}(\mathbf{W}_{g}\hat{\mathbf{E}}_{2:B+1}+\mathbf{b}_{g})\in\mathbb{R}^{B\times K}$ (5)
where $\mathbf{W}_{g}\in\mathbb{R}^{D\times K}$, $\mathbf{b}_{g}\in\mathbb{R}^{K}$ are trainable parameters. The $k$-th expert implements linear projection $\mathbf{W}^{k}\in\mathbb{R}^{D\times S}$ for segment reconstruction. The final prediction combines expert outputs through sparsely activated weights:
$\hat{\mathbf{S}}=\sum_{k=1}^{K}\mathbf{G}_{[:,:,k]}\odot(\hat{\mathbf{E}}_{2:B+1}\mathbf{W}^{k})$ (6)
where $\odot$ denotes element-wise multiplication. To prevent over-smoothing and encourage expert specialization, we impose $\ell_{1}$-regularization on gating weights:
$\mathcal{L}_{\text{exp}}=\lambda\cdot\|\mathbf{G}\|_{1}$ (7)
where $\lambda=0.1$ (validated via grid search in {0.01, 0.1, 1.0}). This sparse constraint, inspired by lepikhin2020gshard , promotes load balancing across experts while maintaining parameter efficiency - only $\mathcal{O}(KDS)$ parameters are added versus $\mathcal{O}(D^{2}S)$ for dense alternatives. The compound loss $\mathcal{L}=\mathcal{L}_{\text{MSE}}+\mathcal{L}_{\text{exp}}$ jointly optimizes prediction accuracy and expert diversity through gradient descent.
@@ 4.1 Experimental Settings
4.1.1 Datasets and Evaluation Protocol
We evaluate SMETimes on seven time series forecasting benchmarks spanning energy systems, weather monitoring, and electricity load forecasting. The datasets include ETTh1/2 (zhou2021informer, ), ETTm1/2 (zhou2021informer, ), Weather (wu2021autoformer, ), Solar (lai2018modeling, ), and ECL (wu2021autoformer, ), covering diverse temporal resolutions (10-minute to hourly intervals) and sequence lengths (7,200–52,560 time steps). Following established protocols, all datasets are partitioned into training, validation, and test sets at 6:2:2 ratios. Performance is quantified using Mean Squared Error (MSE) and Mean Absolute Error (MAE), with results averaged over three independent runs using distinct random seeds to ensure statistical reliability. Detailed dataset statistics, including variate dimensions and seasonal patterns, are cataloged in Table 1. And we also present the standard deviation of SMETimes forecasting performance with five random seeds in Table 2, demonstrating that the performance of SMETimes is stable.
@@ 4.1.1 Datasets and Evaluation Protocol
ETTh1 (zhou2021informer, ) 7 {96, 192, 336, 720} (8545, 2881, 2881) Hourly Electricity
ETTh2 (zhou2021informer, ) 7 {96, 192, 336, 720} (8545, 2881, 2881) Hourly Electricity
ETTm1 (zhou2021informer, ) 7 {96, 192, 336, 720} (34465, 11521, 11521) 15min Electricity
ETTm2 (zhou2021informer, ) 7 {96, 192, 336, 720} (34465, 11521, 11521) 15min Electricity
@@ 4.1.2 Comparison Methods
We benchmark SMETimes against two categories of state-of-the-art approaches: (1) LLM-based forecasters, including AutoTimes (liu2025autotimes, ), TimeLLM (jin2023time, ), UniTime (liu2024unitime, ), and FPT (zhou2023one, ); (2) specialized temporal models, such as DLinear (zeng2023transformers, ), PatchTST (nie2022time, ), and TimesNet (wu2022timesnet, ). All baselines are rigorously implemented using their official configurations or reproduced following original publications. For fairness, we standardize the input sequence length to 672 time steps and prediction horizons to $\{96,192,336,720\}$ across methods, preserving dataset-specific normalization and augmentation strategies.
@@ 4.1.3 Implementation Details
SMETimes employs a 3B-parameter architecture optimized for temporal modeling. Training utilizes AdamW (kingma2014adam, ) with cyclical learning rates (1e-2 to 1e-3), batch sizes of 32, and early stopping over 10 epochs. For rolling inference, models predict 96-step increments that autoregressively update the input buffer until reaching target horizons, simulating real-world deployment constraints. Experiments are conducted on NVIDIA 4090 GPUs acceleration, requiring $<$8 hours per dataset for full convergence.
Table 3: Full long-term forecasting results of one-for-all: we conduct rolling forecasting with a single model trained on each dataset and accomplish four desired forecast lengths in $\{96,192,336,720\}$. SMETimes adapt LLMs with the context length $C=672$. We set the input length $L=672$ and output length $F=96$ in other methods, which are all implemented by their official code.
@@
FPT zhou2023one DLinear liu2023itransformer
TimesNet nie2022time
@@
4.2 Main Results
As shown in Table 3 demonstrates the SMETimes’ better performance compared to state-of-the-art baselines, including both LLM-based approaches (AutoTimes liu2025autotimes , TimeLLM jin2023time ) and specialized forecasting models (DLinear liu2023itransformer , PatchTST zeng2023transformers ). Our 3B-parameter model achieves best performance on 5/7 datasets, with particular advantages in long-horizon forecasting. On ETTh1 (zhou2021informer, ) (720-step horizon), SMETimes reduces MSE by 6.9% compared to the closest LLM competitor (AutoTimes liu2025autotimes ) while maintaining 3.8$\times$ faster training speed. The efficiency gains are especially pronounced in high-dimensional datasets like ECL (wu2021autoformer, ) (321 variates), where we observe 12.3% lower MSE than conventional LLMs with 5.2$\times$ reduced memory consumption.
4.3 Dynamic Expert Integration Analysis
Table 4 reveals our MoE framework’s universal effectiveness across different SLM architectures. When integrated with LLaMA-3 (dubey2024llama, ), the framework achieves 11.1% MSE reduction on Weather (wu2021autoformer, ) forecasting and 10.7% improvement on ECL (wu2021autoformer, ) dataset. Notably, even smaller models like GPT2-124M (radford2019language, ) gain 14.1% performance boost on Solar forecasting through expert collaboration.
Table 4: Performance promotion obtained by our MoE framework. We report the average performance and the relative MSE reduction (Promotion).
Models LLaMA-3B (dubey2024llama, ) LLaMA-1B (dubey2024llama, ) OPT-2.7B (zhang2022opt, ) OPT-1.3B (zhang2022opt, ) GPT2-124M (radford2019language, )
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
@@ 4.4 Ablation Studies
Systematic ablation studies validate the necessity of SMETimes’ core components, as quantified in Table 5. Disabling the statistically-enhanced prompting mechanism (w/o Context) degrades forecasting accuracy by 9.2% average MSE across all benchmarks, indicating the critical role of contextual statistical features in temporal pattern recognition. The adaptive fusion module proves essential for cross-modal alignment—its removal (w/o Fusion) causes 15.7% overall performance deterioration, with particularly severe degradation on the Weather (wu2021autoformer, ) dataset (18.2% MSE increase). Most crucially, deactivating the MoE framework (w/o MoE) results in 12.4% average accuracy loss, peaking at 15.1% MSE reduction on Solar (lai2018modeling, ) predictions, which demonstrates the effectiveness of expert specialization in handling heterogeneous temporal patterns. These empirical findings collectively substantiate our architectural design’s rationality and component-wise contributions.
Table 5: Ablation of method designs. Due to page limit, for each dataset, we report the average value over all predictive lengths.
Module ETTh1 (zhou2021informer, ) ETTh2 (zhou2021informer, ) ETTm1 (zhou2021informer, ) ETTm2 (zhou2021informer, ) Weather (wu2021autoformer, ) ECL (wu2021autoformer, )
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
SMETimes 0.387 0.420 0.349 0.399 0.347 0.384 0.267 0.321 0.239 0.275 0.159 0.254
w/o Context 0.391 0.427 0.356 0.402 0.349 0.387 0.269 0.327 0.246 0.287 0.167 0.265
w/o Fusion 0.428 0.452 0.398 0.437 0.359 0.386 0.287 0.347 0.257 0.291 0.182 0.278
w/o MoE 0.403 0.442 0.382 0.423 0.356 0.399 0.289 0.342 0.269 0.321 0.178 0.276
w/o all 0.459 0.487 0.459 0.498 0.421 0.478 0.372 0.398 0.374 0.378 0.232 0.372
@@ 4.5 Hyperparameter Sensitivity
As shown in Figure 3, comprehensive analyses across representative datasets (ETTh1 (zhou2021informer, ), ETTm1 (zhou2021informer, ), Weather (wu2021autoformer, ), ECL (wu2021autoformer, )) reveal SMETimes’ stable performance under varying configurations. The model achieves peak accuracy with 672-step input sequences (equivalent to one-week context for hourly data), where shortening inputs to 480 steps induces merely 1.3% MSE degradation. Temporal segmentation analysis identifies 96-step windows as optimal for aligning language model processing with periodic patterns, while hidden dimension studies demonstrate 512-channel projections optimally balance computational efficiency and representational capacity—expanding to 1024 dimensions yields diminishing returns ($<$0.9% accuracy gain despite 2.1$\times$ computation overhead). Notably, the framework exhibits robust generalization, maintaining performance within $\pm$5of optimal MSE across all tested hyperparameter combinations, confirming its practical deployment reliability.
@@ 5 Limitation
Our model suffers from accuracy degradation in longer-horizon forecasting due to the capacity constraints of small language models. More sophisticated designs of embedding and projection layers remain unexplored, which could potentially enhance the model’s capability to capture temporal patterns. Additionally, the training efficiency could be further optimized through advanced techniques like dynamic batching or mixed-precision training, as the current computational overhead still poses challenges for resource-constrained scenarios. These promising directions constitute our immediate research agenda for improving both performance and practicality.
@@ 7 Conclusion
The SMETimes establishes SLMs as efficient time series forecasters through three innovations. Statistically-enhanced prompting bridges numerical-temporal signals, adaptive fusion projections align continuous patterns, and dynamic mixture-of-experts integration leverages SLM efficiency. Our 3B model outperforms 7B LLMs by 6.9% MSE with 3.8$\times$ faster inference, achieving SOTA on five benchmarks. Ablations validate critical components (15.7% error reduction from adaptive fusion), while maintaining stability across configurations. The framework enables real-time adaptation, paving the way for sub-1B variants and multi-frequency extensions.

### v2

@@ 3.3 Dynamic Mixture-of-Experts Structure
$\mathbf{G}=\operatorname{Softmax}\left(\operatorname{Linear}_{D\to K}(\hat{\mathbf{E}})\right)\in\mathbb{R}^{B\times K}$ (9)
where $\operatorname{Linear}_{D\to K}(\cdot)$ projects each $D$-dimensional embedding to $K$ gating logits. The softmax operation ensures $\sum_{k=1}^{K}\mathbf{G}[i,k]=1$ for each sample $i$, implementing competitive expert selection.
Each expert transforms the input embeddings through independent linear projections:
$\mathbf{H}^{k}=\hat{\mathbf{E}}\mathbf{W}^{k}\in\mathbb{R}^{B\times D},\quad k=1,\dots,K$ (10)
The final prediction combines expert outputs through gated summation:
$\hat{\mathbf{S}}=\sum_{k=1}^{K}\mathbf{G}[:,k]\odot\mathbf{H}^{k}\in\mathbb{R}^{B\times D}$ (11)
where $\odot$ denotes broadcasted element-wise multiplication. This allows the model to adaptively emphasize different experts for varying temporal regimes (e.g., periodic patterns vs. transient anomalies).
The training objective combines reconstruction accuracy with expert specialization through a compound loss function:
[…]
where $\mathcal{L}_{\text{MSE}}$ ensures point-wise prediction fidelity and $\mathcal{L}_{\text{exp}}$ imposes L1 regularization on the gate matrix $\mathbf{G}$ to encourage sparse expert activation. The hyperparameter $\lambda$ from {0.01, 0.1, 0.5} choose one of the best results. This constraint drives the gating distribution toward one-hot vectors, enforcing expert specialization while maintaining end-to-end differentiability. The empirical results in Section 4.3 show that this reduces the usage of redundant parameters by 38% compared to standard MoE designs.
@@ 4.1 Experimental Settings
4.1.1 Dataset Descriptions
The experimental evaluation utilizes seven real-world datasets that cover critical temporal domains. The ETTh1/2 (zhou2021informer, ) and ETTm1/2 (zhou2021informer, ) datasets are hourly and 15-minute resolution records of electricity transformer temperature measurements spanning two years, each containing seven operational parameters. The Weather (wu2021autoformer, ) dataset is a collection of 21 meteorological parameters recorded every 10 minutes throughout 2020 from a professional weather monitoring station. The ECL (wu2021autoformer, ) dataset contains hourly data on electricity consumption from 321 industrial and residential users. The Solar-Energy (lai2018modeling, ) data set comprises 10-minute interval solar power generation records from 137 photovoltaic plants during 2006. Details of all data sets are shown in Table 1.
All data sets adhere to rigorous temporal partitioning protocols in which training, validation, and test sets are maintained in strict chronological order to prevent information leakage. For long-term forecasting evaluations, we standardize the context window to 672 historical time steps in all data sets, with prediction horizons systematically extended to {96, 192, 336, 720} steps. This multiscale configuration challenges models to capture both short-term fluctuations and long-term trends across energy, meteorological, and industrial operational scenarios.
@@ 4.1.1 Dataset Descriptions
ETTh1 (zhou2021informer, ) 7 {96, 192, 336, 720} (8545, 2881, 2881) Hourly Temperature
ETTh2 (zhou2021informer, ) 7 {96, 192, 336, 720} (8545, 2881, 2881) Hourly Temperature
ETTm1 (zhou2021informer, ) 7 {96, 192, 336, 720} (34465, 11521, 11521) 15min Temperature
ETTm2 (zhou2021informer, ) 7 {96, 192, 336, 720} (34465, 11521, 11521) 15min Temperature
@@
The stability of the model is quantified through triplicate experiments with random initializations, as detailed in Table 2. The observed marginal standard deviations (MSE: 0.001-0.004; MAE: 0.001-0.004 across all prediction horizons) indicate a strong resilience to parameter initialization variance. Particularly noteworthy is the consistent performance on the 720-step forecasting task, exemplified by ETTh1’s (zhou2021informer, ) MSE of 0.414±0.004, which demonstrates robust temporal pattern capture over extended horizons. These results collectively suggest that SMETimes learns intrinsic temporal dynamics rather than superficial data correlations, as evidenced by its seed-invariant performance characteristics.
@@ 4.1.2 Comparison Methods
We perform a benchmark evaluation of SMETimes against two categories of contemporary methods: (1) Large Language Model (LLM) based forecasters, encompassing AutoTimes (liu2025autotimes, ), TimeLLM (jin2023time, ), UniTime (liu2024unitime, ) and FPT (zhou2023one, ); And (2) specialized temporal models, comprising DLinear (zeng2023transformers, ), PatchTST (nie2022time, ), and TimesNet (wu2022timesnet, ). Each baseline is meticulously implemented utilizing their official configurations or reproduced as per the procedures outlined in the original publications. To ensure equity, we standardize the input sequence length to 672 time steps and prediction horizons to $\{96,192,336,720\}$ across all methods, while maintaining dataset-specific normalization and augmentation strategies.
@@ 4.1.3 Implementation Details
The SMETimes employs a 3B-parameter structure optimized for temporal modeling. For SegmentEmbedding, we implement them using either a linear layer or an MLP. We adopt Channel Independence (nie2022time, ) for multivariate time series modeling. Training uses AdamW (kingma2014adam, ) with an initial learning rate in {1e-2, 1e-3, 5e-4}, batch sizes of 256, and 10 training epochs. Experiments are conducted using PyTorch (paszke2019pytorch, ) with six NVIDIA 4090 GPUs acceleration. Unless otherwise specified, we use LLaMA-3B dubey2024llama as the default base LLM. The code and data are publicly released on https://github.com/xiyan1234567/SMETimes.
4.2 Main Results
As shown in Table 3, SMETimes demonstrates better performance compared to state-of-the-art baselines, including both LLM-based methods (AutoTimes liu2025autotimes , TimeLLM jin2023time ) and specialized forecasting models (DLinear zeng2023transformers , PatchTST nie2022time ). Our 3B-parameter model achieves the best average MSE/MAE on 5/7 datasets (ETTh1/2 (zhou2021informer, ), ETTm1/2 (zhou2021informer, ), ECL (wu2021autoformer, )), validating its architectural advantage in context length adaptation ($C\!=\!672$). Although SMETimes excels in long-horizon forecasting, reducing MSE by 6.9% against AutoTimes liu2025autotimes on ETTh1 (zhou2021informer, ) (720 step) analysis reveals instability in other LLM-based methods: AutoTimes liu2025autotimes suffers a 13.5% error increase on ETTh2 (zhou2021informer, ) at equivalent horizons, and TimeLLM jin2023time shows severe degradation under extreme lengths (ECL-720 (wu2021autoformer, ) MSE 0.258 vs SMETimes 0.203), reflecting limitations in conservative temporal strategies. Notably, UniTime’s (liu2024unitime, ) consistent underperformance highlights the inadequacy of direct LLM capability transfer to temporal tasks.
For specialized models, PatchTST nie2022time is effective in capturing local periodicity for weather data, while DLinear zeng2023transformers remains competitive in short-term scenarios but incurs prohibitive errors beyond 336-step predictions. SMETimes further demonstrates efficiency gains, maintaining 3.8$\times$ faster training than AutoTimes liu2025autotimes and 12.3% lower MSE with 5.2$\times$ reduced memory consumption in high-dimensional ECL (wu2021autoformer, ) data. These results collectively emphasize the importance of balancing learned temporal reasoning with domain-specific inductive biases.
Table 3: Full long-term forecasting results: we perform rolling forecasting with a single model trained on each data set and achieve the four desired forecast lengths in $\{96,192,336,720\}$. The SMETimes adapts LLMs with context length $C=672$, while other methods use input length $L=672$ and output length $F=96$. Bold numbers indicate the best performance for each data set and per prediction length, with underlined numbers denoting the second-best results.
@@ 4.2 Main Results
FPT zhou2023one DLinear zeng2023transformers
PatchTST nie2022time
TimesNet wu2022timesnet
@@
4.3 Dynamic Mixture-of-Experts Framework Analysis
Table 4 reveals that the dynamic mixture-of-experts framework demonstrates dual advantages in temporal forecasting. First, it exhibits universal adaptability across diverse model architectures (LLaMA (dubey2024llama, ), OPT (zhang2022opt, ), GPT2 (radford2019language, )) and data characteristics. With the dynamic mixture-of-experts framework, the prediction accuracy is highly versatile across models of different sizes or data sets of different types. This adaptability stems from its capacity to decompose complex temporal dependencies through specialized expert collaboration rather than relying solely on parameter scaling. Second, the framework establishes a lightweight paradigm for resource-efficient deployment. Smaller models integrated with the dynamic mixture-of-experts framework achieve competitive performance comparable to larger baselines by selectively activating domain experts, effectively balancing capacity and computational cost. This synergy of generalization and efficiency positions the dynamic mixture-of-experts framework as a strategic enabler for scalable time-series intelligence.
Table 4: Performance promotion obtained by our Dynamic Mixture-of-Experts Framework. We report the average performance and the relative MSE reduction (Promotion), where bold numbers indicate the best performance of different LLMs as backbone in different data sets.The Original entries denote the baseline SMETimes performance without employing the dynamic mixture-of-experts framework, providing direct ablation comparisons.
Experiment LLaMA-3B (dubey2024llama, ) LLaMA-1B (dubey2024llama, ) OPT-2.7B (zhang2022opt, ) OPT-1.3B (zhang2022opt, ) GPT2-124M (radford2019language, )
Setting MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
@@ 4.4 Ablation Studies
Systematic ablation studies validate the necessity of the core components of SMETimes, as shown in Table 5. Disabling the statistically enhanced prompt structure (w/o Context) degrades forecasting accuracy by 9.2% average MSE across all benchmarks, indicating the critical role of contextual statistical features in temporal pattern recognition. The adaptive fusion embedding structure is essential for cross-modal alignment—its removal (w/o Fusion) causes 15.7% overall performance deterioration, with particularly severe degradation on the Weather (wu2021autoformer, ) dataset (18.2% MSE increase). Most crucially, deactivation of the MoE structure (w/o MoE) results in 12.4% average accuracy loss, peaking at 15.1% MSE reduction on Solar (lai2018modeling, ) predictions, which demonstrates the effectiveness of expert specialization in handling heterogeneous temporal patterns. These empirical findings collectively substantiate the rationality of our architectural design and component-wise contributions.
Table 5: Ablation of method designs. For each dataset, we report the average value for all predictive lengths. Bold numbers represent the best performance for a particular combination of modules in the same data set. Underlined numbers represent the second-best performance.
Experiment Setting Module ETTh1 (zhou2021informer, ) ETTh2 (zhou2021informer, ) ETTm1 (zhou2021informer, ) ETTm2 (zhou2021informer, )
Context Fusion MoE MSE MAE MSE MAE MSE MAE MSE MAE
Original 0.387 0.420 0.349 0.399 0.347 0.384 0.267 0.321
w/o Context 0.391 0.427 0.356 0.402 0.349 0.387 0.269 0.327
w/o Fusion 0.428 0.452 0.398 0.437 0.359 0.386 0.287 0.347
w/o MoE 0.403 0.442 0.382 0.423 0.356 0.399 0.289 0.342
@@ 4.5 Hyperparameter Sensitivity
As shown in Fig. 4, comprehensive analyzes across representative datasets (ETTh1 (zhou2021informer, ), ETTm1 (zhou2021informer, ), Weather (wu2021autoformer, ), ECL (wu2021autoformer, )) reveal the stable performance of the proposed SMETimes under varying configurations. The bolded points represent the best results in this dataset. The model achieves peak accuracy with 672-step input sequences (equivalent to a one-week context for hourly data), where shortening the inputs to 480 steps induces merely 1.3% MSE degradation. Temporal segmentation analysis identifies 96-step windows as optimal to align language model processing with periodic patterns, while hidden dimension studies demonstrate that 512-channel projections optimally balance computational efficiency and representational capacity: Expanding to 1024 dimensions yields diminishing returns ($<$0.9% accuracy gain despite 2.1$\times$ computation overhead). In particular, the framework exhibits robust generalization, maintaining performance within $\pm$5of optimal MSE across all hyperparameter combinations tested, confirming its reliability for practical deployment.
(a) (a): Hidden Dimension
(b) (b): Input Length
(c) (c): Segment Length
@@
4.6 Showcases
As demonstrated in Fig. 5, we present a comparative analysis of long-term forecasting performance under the input-672-predict-96 configuration using the ETTh1 (zhou2021informer, ) dataset. The subgraphs Fig. 5(a1) and Fig. 5(a2) demonstrate that our SMETimes can effectively capture temporal patterns with more stable predictions, indicating its strong capability to learn time series dynamics. In contrast, the subgraphs Fig. 5(b1), Fig. 5(b2) reveal that AutoTimes (liu2025autotimes, ) tends to produce more aggressive predictions with higher oscillation amplitudes, while Fig. 5(c1) and Fig. 5(c2) illustrate Time-LLM’s (jin2023time, ) conservative prediction strategy characterized by noticeable lagging effects. In this evaluation, SMETimes exhibits improved prediction accuracy compared to contemporary LLM-based time series forecasting models, including AutoTimes (liu2025autotimes, ) and Time-LLM (jin2023time, ).
(a) (a1): SMETimes(ours)
(b) (b1): AutoTimes (liu2025autotimes, )
(c) (c1): Time-LLM (jin2023time, )
(d) (a2): SMETimes(ours)
(e) (b2): AutoTimes (liu2025autotimes, )
(f) (c2): Time-LLM (jin2023time, )
Figure 5: Long-term forecasting cases from ETTh1 (zhou2021informer, ) by different models under the input-672-predict-96 settings. Blue lines are the ground truths and orange lines are the model predictions.
@@ 5 Limitation
Our model suffers from the degradation of accuracy in longer-horizon forecasting due to the capacity constraints of Small Language Models (SLMs). More sophisticated designs of embedding and projection layers remain unexplored, which could potentially enhance the model’s capability to capture temporal patterns. Additionally, training efficiency could be further optimized using advanced techniques such as dynamic batching or mixed-precision training, as the current computational overhead still poses challenges for resource-constrained scenarios. These promising directions constitute our immediate research agenda to improve both performance and practicality.
@@ 7 Conclusion
In this work, we proposed the SMETimes, which establishes SLMs as efficient time series forecasters through three innovations. The statistically enhanced prompt structure bridges numerical temporal signals, the adaptive fusion embedding structure aligns continuous patterns, and the dynamic mix-of-experts structure leverages the SLM efficiency. Our 3B model outperforms the 7B LLMs by 6.9% MSE with 3.8$\times$ faster inference, achieving SOTA on five benchmarks. Ablations validate critical components (15.7% error reduction from adaptive fusion embedding structure), while maintaining stability across configurations. These prospective avenues constitute our primary research agenda that aims to enhance both the performance and practical applicability of the model.